OpenStudy (anonymous):

1. The distance traveled by an object can be modeled by the equation d = ut + 0.5at2 where d = distance, u = initial velocity, t = time, and a = acceleration. Solve this formula for a. Show all steps in your work.\

OpenStudy (the_fizicx99):

\(\ \sf d = ut + \dfrac{1}{2}at^2 \) What they mean by "Solve this formula for a" is to isolate a such that "a" = something Start by doing the inverse of everything, subtract ut from both sides; \(\ \sf d - ut = \cancel{ut} - \cancel{ut} + \dfrac{1}{2}at^2 \implies d - ut = \dfrac{1}{2}at^2\) Any ideas on how to continue on your own?

OpenStudy (anonymous):

multiply 1/2 by each side?

OpenStudy (the_fizicx99):

No, keep in mind you're already multiplying. so you would do the inverse of that, which is division

OpenStudy (the_fizicx99):

1/2 = 0.5 d - ut = 0.5at^2, does that help?

OpenStudy (anonymous):

okay ,so divide that by each side?

OpenStudy (the_fizicx99):

Yeah.  Therefore, 2(d - ut) = at^2. Solve for "a" by doing the inverse and dividing.